# On Vanishing Gradients, Over-Smoothing, and Over-Squashing in GNNs: Bridging Recurrent and Graph Learning

### v1

@@ 3.3 GNN-SSM: Improving the training dynamics of GNNs through state-space models
To allow direct control of the signal propagation dynamics of any GNN, we can rewrite its layer-to-layer update as a state-space model. Concretely, we express the update as
$\mathbf{H}^{(k+1)}$ $=\mathbf{\Lambda}\mathbf{H}^{(k)}+\mathbf{B}\mathbf{X}^{(k)}$
$=\mathbf{\Lambda}\mathbf{H}^{(k)}+\mathbf{B}\mathbf{F}_{\boldsymbol{\theta}}(\mathbf{H}^{(k)},k),$ (9)
where we refer to $\mathbf{\Lambda}$ as the state transition matrix and $\mathbf{B}$ as the input matrix, [^2] Here, we deviate from the traditional state-space formalism, which uses $\mathbf{A}$ as the state transition matrix, since we use this notation for the adjacency. and $\mathbf{F}_{\boldsymbol{\theta}}(\mathbf{H}^{(k)},k)$ as a time-varying coupling function which connects each node to some neighborhood. We refer to the model defined in Equation (9) as GNN-SSM. From an RNN perspective, $\mathbf{\Lambda}$ plays the role of the “memory”, in charge of recalling all the representations at each layer at the readout layer, while the neighborhood aggregation plays the role of an input injected into the state via $\mathbf{B}$. In traditional GNNs, this recurrent memory mechanism is absent, so these models act in a memoryless way: features at one layer do not explicitly store or retrieve past information in the way a stateful model would.
@@
The result above shows that the spectrum of the Jacobian is controlled through the eigenvalues of the memory matrix $\mathbf{\Lambda}$. Since $\text{eig}(\Gamma)\approx 0$ (see previous subsection), it suffices to have $\text{eig}(\Lambda)\approx 1$ to bring the vectorized Jacobian to the edge of chaos. We empirically validate this in Figure 3.
Figure 3: Vectorized Jacobian for different models. Left: GCN. Right: GCN-SMM with $\text{eig}(\mathbf{\Lambda})\approx 1$ and $\text{eig}(\mathbf{B})\approx 0.1$.
@@
For simplicity and clarity of conclusions, we consider $\mathbf{\Lambda}$ and $\mathbf{B}$ to be shared across layers and fixed (i.e., not trained by gradient descent). Only the coupling function $\mathbf{F}_{\boldsymbol{\theta}}$ is optimized. Empirically, we observe that this simpler scheme actually improves downstream performance in some settings. We highlight, however, that this is the most simple instance of a more general framework that aims to incorporate ideas from recurrent processing into GNNs without losing permutation-equivariance. One could easily extend this state-space idea to include more complex gating (Hochreiter & Schmidhuber, 1997; Rusch et al., 2023b) or other constraints on the state transition matrix.
@@
Message of the Section: GNNs, especially GCNs and GATs, experience a phenomenon we term “extreme gradient vanishing” due to the use of a normalized adjacency for the message passing operation. Reinterpreting any GNN as a recurrent model enables direct control of the Jacobian spectrum, which mitigates this issue while preserving architectural generality.
4 How is Over-smoothing linked to Vanishing Gradients?
@@
In this section, we study the practical implications of the mechanism causing vanishing gradients in GNNs in relation to the over-smoothing phenomenon. We show theoretically how over-smoothing is a consequence of GNN layers acting as contractions, which make node features collapse to a zero fixed point under certain conditions. We experimentally validate our points by analyzing Dirichlet energy, node feature norms, and node classification performance for increasing numbers of layers. Overall, we believe this section provides a more practical and general understanding of over-smoothing, by analyzing any GNN from the point of view of its layer-wise Jacobians.
4.1 Over-smoothing secretly occurs due to contractions in the Jacobian
Over-smoothing describes the tendency of node features to become too similar to each other as more layers are added in GNNs (Cai & Wang, 2020; Rusch et al., 2023a). This phenomenon is largely due to the nature of the operations that are performed by GNN layers and their relationship with heat equations over graphs. Consequently, a common way of measuring over-smoothing in GNNs is via the graph Dirichlet energy $\mathcal{E}(\mathbf{H})$. Given a graph node signal $\mathbf{H}\in\mathbb{R}^{n\times d}$ on a graph $G$, $\mathcal{E}(\mathbf{H})$ takes the form:
$\mathcal{E}(\mathbf{H})=\text{tr}\left(\mathbf{H}^{\top}\boldsymbol{\Delta}\mathbf{H}\right)=\sum_{(u,v)\in E}\left\lVert\mathbf{h}_{u}-\mathbf{h}_{v}\right\rVert^{2},$ (11)
where $\boldsymbol{\Delta}$ is the (unweighted) graph Laplacian (Chung, 1997). The graph Dirichlet energy measures the smoothness of a signal over a graph and will be minimized when the signal is constant over each node – at least when using the non-normalized Laplacian. Notably, models like GCNs can be seen as minimizers of the Dirichlet energy [^3] In general, this depends on the spectral properties of the learned weight matrices and on the details of the non-linearities used (Di Giovanni et al., 2022). (Bodnar et al., 2022; Di Giovanni et al., 2022).
@@ 4.1 Over-smoothing secretly occurs due to contractions in the Jacobian
Figure 4: Experimental evaluation on Cora for an increasing number of layers. Left: Dirichlet Energy evolution for different $||\Lambda||_{2}$. Middle: 2-Dimensional feature projection evolution with a fixed point at zero. Right: Node classification performance.
We consider in our analysis GNN layers as in Equation 3. We view a GNN layer as a map $f_{k}:\mathbb{R}^{nd}\to\mathbb{R}^{nd}$ and construct a deep GNN $f$ via composition of $K$ layers, i.e. $f=f_{K}\circ\dots\circ f_{1}$. We also require the condition of our non-linearity $\sigma$ that $\sigma(0)=0$, noting that this is the case for ReLU and $\tanh$, for example. We start by showing that in such a model $\mathbf{0}$ is a fixed point. [^4] In our analysis, it is important that the input to the GNN is a vector in $\mathbb{R}^{nd}$ rather than a matrix in $\mathbb{R}^{n\times d}$, as the Jacobians and norms are different for the two cases. For this reason, it is important to take care in the definitions of these objects.
@@
Let $\mathbf{J}_{f}\in\mathbb{R}^{nd\times nd}$ denote the layer-wise Jacobian of a GNN $f$. The supremum of the Jacobian (if well-defined) of $f$ over a convex set $U$ corresponds to the Lipschitz constant $\left\lVert{f}\right\rVert_{\text{Lip}}$ (Hassan et al., 2002), i.e.
$\left\lVert{f}\right\rVert_{\text{Lip}}=\sup_{\mathbf{H}\in U}\left\lVert\mathbf{J}_{f}(\mathbf{H})\right\rVert,$ (12)
where by submultiplicativity of Lipschitz constants we have that $\left\lVert{f}\right\rVert_{\text{Lip}}\leq\prod_{k=1}^{K}\left\lVert{f_{k}}\right\rVert_{\text{Lip}}$. In the case of ReLU, as the function is not differentiable at 00, one has to consider the (Clarke) generalized Jacobian (Jordan & Dimakis, 2020), a detail that we ignore to ease notation as it is not important for the scope of our work. We point to the Appendix A.2 for a more detailed explanation of the objects in question. A Lipschitz function $f$ is contractive if $\left\lVert{f}\right\rVert_{\text{Lip}}<1$. We now assume that $\left\lVert{f_{k}}\right\rVert_{\text{Lip}}<1$ for all $k$, meaning that each layer is a contraction mapping. [^5] Note that the analysis holds for any submultiplicative matrix norm. We derive the following result:
@@
In other words, if layers $f_{k}$ are contractive, their repeated application will monotonically converge to the unique fixed point $\mathbf{0}$, by Lemma 4.1 and the Banach fixed point theorem. Each GNN layer has therefore the effect of contracting the input vectors, reducing their norms. The rate of reduction is dictated by the spectral norm of the Jacobian of that layer and therefore by the spectral norm of $\mathbf{W}_{k}$, as shown in Section 3. In Proposition 4.3, we show how the layer-wise Jacobians relate to the Dirichlet energy.
@@
This result shows that the energy is directly controlled by the norm of the input signal $\mathbf{H}$ and by the contracting effect of the layers $f_{k}$. The repeated application of contractive layers results in the Dirichlet energy being artificially lowered as the signals are gradually reducing in norm.
This provides a much more practical explanation for over-smoothing than the one usually encountered in the literature. Contrary to common consensus, which explains over-smoothing by showing that the signal is projected into the $1$-dimensional kernel of the graph Laplacian. We instead describe over-smoothing as occurring due to the contractive nature of GNNs and their inputs converging in norm to exactly $\mathbf{0}$.
@@
The most important takeaway of the analysis above is that vanishing gradients are directly connected to over-smoothing. In particular, the same mechanism that causes gradient vanishing issues - the non-isometric propagation of signals due to contractions of the Jacobian - is responsible for the collapse of all features to a unique zero fixed point where the Dirichlet energy is minimized. The quick collapse of traditional graph convolutional and attentional models can therefore be understood from the extreme gradient vanishing result introduced in Section 3.
@@
This result also provides a connection between the study of GNNs and the study of signal propagation (or dynamical isometry) in feedforward networks (Saxe et al., 2013; Poole et al., 2016; Pennington et al., 2017) and recurrent neural networks (Hochreiter & Schmidhuber, 1997; Arjovsky et al., 2016; Orvieto et al., 2023). In the dynamical isometry literature, the primary interest is to improve the learning times of deep feedforward networks, whereas the recurrent neural network literature is interested in memory and long-range information retrieval. We highlight that translating ideas from these fields of study will enable the construction of deep and performant GNNs, even though these techniques were originally developed with other objectives in mind. This also serves as an explanation of why simple modifications such as residual connections or normalization worked in practice to mitigate over-smoothing, given their links to dynamical isometry (Yang et al., 2019; Meterez et al., 2024).
@@
Finally, we highlight that this result provides an objective evaluation metric to gauge whether a GNN will over-smooth or not. We hope that the eigenalysis of the Jacobian will become a widespread empirical test used for this purpose. In Appendix D.2, we showcase how several models with edge-of-chaos Jacobians result in no over-smoothing.
4.2 Experimental validation of theoretical results
@@
To validate the theory above, we perform a series of empirical tests. In particular, we check the evolution of the Dirichlet energy, latent vector norms, and node classification accuracy on the Cora dataset as the number of layers of different models is increased. The results are presented in Figure 4. Further, we present additional experiments for different graph structures and models in Appendix D.2.
@@
From Figure 4, we see that one can exactly control the evolution of the Dirichlet energy of the system through the spectrum of the Jacobian, which can, in turn, be modified through the spectrum of $\mathbf{\Lambda}$. Furthermore, this shrinks faster the lower the norm of the Jacobian is, which validates Proposition 4.3. From the phase plot describing the two-dimensional evolution of the features, it is also clear that these converge to a unique fixed point at zero, in line with Proposition 4.2. Beyond a Dirichlet energy analysis of the system, notice that node classification performance does not deteriorate when $\text{eig}(\mathbf{\Lambda})\approx 1$, and improves over simply applying an SSM layer with no modulation of the hyperparameters or a residual connection. The dynamics of the GNN in this setting are shown in Figure 1.
@@
Message of the Section: There exists a direct link between the over-smoothing phenomenon in GNNs and the appearance of vanishing gradients. In particular, for contractive layerwise Jacobians and certain nonlinearities, node features collapse to a zero fixed point and thus minimize the Dirichlet energy. Hence, analyzing the spectrum of the layer-wise Jacobians will reveal if a GNN will over-smooth or not. Furthermore, borrowing techniques from RNNs to design new GNNs is expected to be an effective strategy to prevent over-smoothing.

### v2

@@ 3.3 GNN-SSM: Improving the training dynamics of GNNs through state-space models
To allow direct control of the signal propagation dynamics of any GNN, we can rewrite its layer-to-layer update as a state-space model. Concretely, we express the update as {eqbox}
$\mathbf{H}^{(k+1)}{}^{\mathsf{T}}$ $=\mathbf{\Lambda}\,\mathbf{H}^{(k)}{}^{\mathsf{T}}+\mathbf{B}\,\mathbf{X}^{(k)}{}^{\mathsf{T}}$
$=\mathbf{\Lambda}\,\mathbf{H}^{(k)}{}^{\mathsf{T}}+\mathbf{B}\,\mathbf{F}_{\boldsymbol{\theta}}\bigl(\mathbf{H}^{(k)},\,k\bigr)^{\mathsf{T}}$ (9)
where we refer to $\mathbf{\Lambda}$ as the state transition matrix and $\mathbf{B}$ as the input matrix, [^3] Here, we deviate from the traditional state-space formalism, which uses $\mathbf{A}$ as the state transition matrix, since we use this notation for the adjacency. Further, we employ transposes to increase the resemblance to the traditional SSM updates. and $\mathbf{F}_{\boldsymbol{\theta}}(\mathbf{H}^{(k)},k)$ as a time-varying coupling function which connects each node to some neighborhood. We refer to the model defined in Equation (9) as GNN-SSM. From an RNN perspective, $\mathbf{\Lambda}$ plays the role of the “memory”, in charge of recalling all the representations at each layer at the readout layer, while the neighborhood aggregation plays the role of an input injected into the state via $\mathbf{B}$. In traditional GNNs, this recurrent memory mechanism is absent, so these models act in a memoryless way: features at one layer do not explicitly store or retrieve past information in the way a stateful model would.
@@
The result above shows that the spectrum of the Jacobian is controlled through the eigenvalues of $\mathbf{\Lambda}$. If we have that the spectrum of $\Gamma$ is around zero, it suffices to have $\text{eig}(\Lambda)\approx 1$ to bring the vectorized Jacobian to the edge of chaos. We empirically validate this in Figure 2.
@@
For simplicity and clarity of conclusions, we consider $\mathbf{\Lambda}$ and $\mathbf{B}$ to be shared across layers and fixed (i.e., not trained by gradient descent) to guarantee the desired properties. To construct $\mathbf{\Lambda}$, we first generate a random unitary matrix, which has all eigenvalues on the unit circle, and then scale it to control the spectral radius [^4] An alternative strategy consists in computing the matrix eigendecomposition, manually place the eigenvalues where desired, and reconstruct the matrix., allowing us to design with precise control over the system dynamics. Only the coupling function $\mathbf{F}_{\boldsymbol{\theta}}$ is optimized. Empirically, we observe that this simpler scheme actually improves downstream performance in some settings. We highlight, however, that this is the most simple instance of a more general framework that aims to incorporate ideas from recurrent processing into GNNs without losing permutation-equivariance. One could easily extend this state-space idea to include more complex gating hochreiter1997long or other constraints on the state transition matrix henaff2016 .
@@
Message of the Section: GCNs and GATs experience a phenomenon we term “extreme gradient vanishing" due to the use of a normalized adjacency for the message passing operation. Reinterpreting any GNN as a recurrent model enables direct control of the Jacobian spectrum, which mitigates this issue while preserving architectural generality.
4 How does Extreme Gradient Vanishing affect Over-smoothing?
@@
In this section, we study the practical implications of the mechanism causing vanishing gradients in GNNs in relation to node feature evolution. We show empirically and theoretically how GNN layers acting as contractions make node features collapse to a fixed point. We experimentally validate our points by analyzing Dirichlet energy, node feature norms, and node classification performance for increasing numbers of layers. Overall, we believe this section provides a more practical and general understanding of the consequences of extreme vanishing in GNNs by analyzing them from the point of view of their layer-wise Jacobians.
4.1 A contractive GNN leads to node feature collapse
We consider in our analysis GNN layers as in Equation 3. We view a GNN layer as a map $f_{k}:\mathbb{R}^{nd}\to\mathbb{R}^{nd}$ and construct a deep GNN $f$ via composition of $K$ layers, i.e. $f=f_{K}\circ\dots\circ f_{1}$. Let $\mathbf{J}_{f}\in\mathbb{R}^{nd\times nd}$ denote the layer-wise Jacobian of a GNN $f$.  [^5] In our analysis, it is important that the input to the GNN is a vector in $\mathbb{R}^{nd}$ rather than a matrix in $\mathbb{R}^{n\times d}$, as the Jacobians and norms are different for the two cases. For this reason, it is important to take care in the definitions of these objects. The supremum of the Jacobian (if well-defined) of $f$ over a convex set $U$ corresponds to the Lipschitz constant $\left\lVert{f}\right\rVert_{\text{Lip}}$ hassan2002nonlinear , i.e. $\left\lVert{f}\right\rVert_{\text{Lip}}=\sup_{\mathbf{H}\in U}\left\lVert\mathbf{J}_{f}(\mathbf{H})\right\rVert$, where by submultiplicativity of Lipschitz constants we have that $\left\lVert{f}\right\rVert_{\text{Lip}}\leq\prod_{k=1}^{K}\left\lVert{f_{k}}\right\rVert_{\text{Lip}}$. We point to the Appendix A.2 for a more detailed explanation of the objects in question. A Lipschitz function $f$ is contractive if $\left\lVert{f}\right\rVert_{\text{Lip}}<1$. We now assume that $\left\lVert{f_{k}}\right\rVert_{\text{Lip}}<1$ for all $k$, meaning that each layer is a contraction mapping. [^6] Note that the analysis holds for any submultiplicative matrix norm.
@@ 4.1 A contractive GNN leads to node feature collapse
Figure 3: Experimental evaluation on Cora for an increasing number of layers. Left: Dirichlet Energy evolution for different $\Lambda||_{2}$. Middle: 2-Dimensional random feature projection evolution with a fixed point at zero. Right: Node classification performance.
Lemma 4.1 (Banach Fixed Point Theorem banach1922 ).
Let $f$ be an operator with Lipschitz constant $\left\lVert{f}\right\rVert_{\text{Lip}}<1$. Then for any starting point $\mathbf{x}_{0}$, the fixed‐point iteration $\mathbf{x}_{n+1}=f(\mathbf{x}_{n})$ converges to the unique fixed point of $f$ at a linear rate $O\bigl(1/\left\lVert{f}\right\rVert_{\text{Lip}}^{n}\bigr)$
From Lemma 4.1 above and the extreme gradient vanishing results presented in Section 3.2, we can distinguish two cases for contractive GNN layers: (1) With shared layers (as in di2022graph ), repeated applications of the GNN will result in convergence to a unique fixed point; (2) with repeated application of non-shared but highly contracting layers, the overall GNN function will converge to, or very close to, a unique fixed point in a single forward pass due to the low Lipschitz constant of the overall GNN. Both cases result in convergence towards a fixed point as all nodes evolve using common transformations. In practice, we observe that node representations tend to collapse to a zero norm node state, see Figure 3. To study this further, we consider a GNN update, under the following assumption, which is consistent with the setup in (3):
@@
Then, we have that node representations will evolve as in Proposition 4.3 presented next.
{eqbox}
@@
In other words, in the setting we have considered, if layers $f_{k}$ are contractive, their repeated application will monotonically converge to the unique fixed point $\mathbf{0}$, by Lemmas 4.2 and 4.1, which serves to explain the behavior observed empirically. We note that this is similar to the analysis in roth2024rank , but we highlight that our analysis is much more broad, as it applied to a number of models beyond GCNs, and is it only required knowledge of behavior of each layer through the Lipschitz constant. Furthermore, we emphasize the important connection between the Lipschitz constant and the vanishing gradients problem, which are linked through the Jacobian.
The collapse of node features to a single point has been typically described as over-smoothing. In particular, over-smoothing describes the tendency of node features to become too similar to each other as more layers are added in GNNs cai2020note . A common way of measuring over-smoothing in GNNs, see rusch2022graph ; rusch2023survey , is via the unnormalized Dirichlet energy $\mathcal{E}(\mathbf{H})$. [^7] We offer a more precise characterization oh how this relates to the normalized Dirichlet energy of node features in Appendix E. Given a feature matrix $\mathbf{H}\in\mathbb{R}^{n\times d}$ on an unweighted graph $G$, $\mathcal{E}(\mathbf{H})$ takes the form:
$\mathcal{E}(\mathbf{H})=\text{tr}\left(\mathbf{H}^{\top}\boldsymbol{\Delta}\mathbf{H}\right)=\sum_{(u,v)\in E}\left\lVert\mathbf{h}_{u}-\mathbf{h}_{v}\right\rVert^{2},\vskip-2.84526pt$ (12)
where $\boldsymbol{\Delta}$ is the unnormalized graph Laplacian chung1997spectral . The Dirichlet energy measures the smoothness of a signal over a graph and will be minimized when the signal is constant over each node – at least when using the unnormalized Laplacian. In Proposition 4.4, we show how the layer-wise Jacobians relate to the unnormalized Dirichlet energy.
{eqbox}
@@
This result shows that the energy is directly controlled by the norm of the input signal $\mathbf{H}$ and by the contracting effect of the layers $f_{k}$. The repeated application of contractive layers results in the unnormalized Dirichlet energy being artificially lowered as signals are gradually reduced in norm.
@@
1) The most important takeaway of the analysis above is that vanishing gradients are directly connected to over-smoothing through the Lipschitz constant. In particular, the same mechanism that causes gradient vanishing issues, is responsible for the collapse of all features to a unique zero fixed point (i.e. zero feature collapse) where the Dirichlet energy is minimized. The quick collapse of traditional graph convolutional and attentional models can also be understood from the extreme gradient vanishing result introduced in Section 3. As such, while much of the literature, starting with oono2020graph ; cai2020note , has attributed over-smoothing to iterative aggregation and convergence to a rank-one subspace, our analysis shows that this phenomenon is instead an artifact of representations collapsing to zero. As later noted by arnaiz2025oversmoothing (without any formal argument), over-smoothing is not necessarily the cause of performance degradation. Here, we provide a theoretical explanation for the poor performance of MPNNs at large depths, linking it to inherent trainability limitations arising from vanishing (and in some cases exploding) gradients.
@@
2) This result also provides a connection between the study of GNNs and the study of signal propagation (or dynamical isometry) in feedforward networks saxe2013exact ; poole2016exponential ; pennington2017resurrecting and recurrent neural networks hochreiter1997long ; arjovsky2016unitary ; orvieto2023resurrecting . In the dynamical isometry literature, the primary interest is to improve the learning times of deep feedforward networks, whereas the recurrent neural network literature is interested in memory and long-range information retrieval. We highlight that connecting ideas from these fields of study will enable the design of new models that benefit from both worlds, even though these techniques were originally developed with other objectives in mind. This also serves as an explanation of why simple modifications such as residual connections or normalization worked in practice to mitigate over-smoothing, given their links to dynamical isometry yang2019mean ; metereztowards .
@@
3) Finally, we highlight that this result provides an objective evaluation metric to gauge whether a GNN will over-smooth or not. We hope that the eigenanalysis of the Jacobian will become a widespread empirical test used for this purpose.
4.2 Empirical validation of theoretical results
@@
To validate the theory above, we perform a series of empirical tests. In particular, we check the evolution of the Dirichlet energy, latent vector norms, and node classification accuracy on the Cora dataset as the number of layers of different models is increased. The results are presented in Figure 3. Further, we present additional experiments for different graph structures and models in Appendix D.1, and showcase how several models with edge-of-chaos Jacobians result in no over-smoothing, reinforcing the generality of this result beyond classical MPNNs.
Figure 4: Left: Evolution of Dirichlet Energy on the Cora dataset for GIN and Gated-GCN. Right: Histograms of eigenvalue spectra of layer-to-layer Jacobians for GIN and Gated-GCN.
@@
From Figure 3, we see that one can exactly control the evolution of the Dirichlet energy of the system through the spectrum of the Jacobian, which can, in turn, be modified through the spectrum of $\mathbf{\Lambda}$ in the GNN-SSM model. Furthermore, this shrinks faster the lower the norm of the Jacobian is, which validates Proposition 4.4. Beyond a Dirichlet energy analysis of the system, notice that node classification performance does not deteriorate when $\text{eig}(\mathbf{\Lambda})\approx 1$, and improves over simply applying an SSM layer with no modulation of the hyperparameters or a residual connection without gating. The dynamics of the GNN in this setting are shown in Figure 1.888Note that in the case of GCN, it is sufficient to scale weight norms to prevent the Dirichlet energy from collapsing to zero with additional layers, as the Dirichlet energy only depends on the maximum eigenvalue of the Jacobian. However, obtaining good performance on node-classification tasks required to transport the whole spectrum to be centred around the edge of chaos to obtain good performance. We elaborate on this in App. E
Results for Additional MPNNs. In this section, we have focused on GCN and GAT because their layer operators admit clean characterizations, which makes the analysis mathematically tractable. However, the arguments extend to any GNN, where it suffices to examine the eigenvalues of the layer-wise Jacobians. To make this concrete, in Figure 4 we plot (i) histograms of the Jacobian eigenvalues for GIN xu2018powerful , and Gated-GCN bresson2017residual , and (ii) the evolution of each model’s Dirichlet energy as the number of layers increases. We observe that GIN exhibits occasional outliers with $|\lambda|>1$ and Gated-GCN displays a sizeable mass above the unit circle, which indicates unstable dynamics. This perspective helps explain several empirical findings reported in arnaiz2025oversmoothing : the divergence of Dirichlet energy in some architectures follows directly from norm expansion driven by unstable Jacobians, whereas the zero feature collapse seen in GCN and GAT reflects their contraction. Consequently, while not all GNNs oversmooth (some instead become unstable and blow up) the generally poor performance of many GNNs at large depths can be understood through their learning dynamics: collapse under contraction (as in GCNs and GATs, especially with ReLUs) or divergence under instability.
@@
Message of the Section: There exists a direct link between the over-smoothing phenomenon in GNNs and the appearance of vanishing gradients. In particular, for contractive layerwise Jacobians and certain nonlinearities, over-smoothing is an artifact of node features collapsing to a zero fixed point which minimizes the Dirichlet energy. Certain GNNs will experience node feature divergence due to the opposite effect of exploding gradients. In both cases, GNNs fail to train at large depth due to issues of trainability. Hence, analyzing the spectrum of the layer-wise Jacobians will reveal if a GNN will over-smooth and train at large depth. Furthermore, borrowing techniques from RNNs to design new GNNs is expected to be an effective strategy to prevent over-smoothing and train deep GNNs.